OpenStudy (anonymous):

Multiply using duplation and mediation. 17 × 48

OpenStudy (anonymous):

799 OKAY I"M I RIGHT!

OpenStudy (anonymous):

I still want to know what duplation and mediation is....

hartnn (hartnn):

and 17x48 is not 799

OpenStudy (anonymous):

Darn -_-

OpenStudy (anonymous):

Can I have the steps

OpenStudy (anonymous):

Should I duplicate or meditate?

hartnn (hartnn):

17X48 =17*2*24 =34*24

hartnn (hartnn):

34*24 =68*12 =136*6 =272*3 =816*1 =816

OpenStudy (anonymous):

Double and half.....

hartnn (hartnn):

or half and double....

OpenStudy (anonymous):

U cheated here 816*1 :-)

OpenStudy (anonymous):

OH OKAY I see

hartnn (hartnn):

:P not 'half' exactly

OpenStudy (anonymous):

"OH OKAY I see" Glad someone does.......

hartnn (hartnn):

its not only halving and doubling..... go on reducing one number......